# T cell receptor binding prediction: A machine learning revolution

### v1

@@ 5 Outlook
The accurate prediction of TCR specificity is crucial for multiple clinical applications such as the development of safer and more efficient immunotherapies. It can also deepen our understanding of autoimmune diseases. Despite the progress, current TCR prediction models face significant challenges due to insufficient and often biased data, notably concerning epitope information. This limitation restricts the models’ capacity to generalize to new epitopes, which is critical, for instance, to predict potential cross-reactive events in T cell-based therapies.
In recent years, we have seen the rapid release and evolution of machine learning models exploiting supervised and unsupervised learning approaches. The advent of protein language models (PLMs), which leverage transfer learning to enhance performance in data-scarce domains, is revolutionizing the field and showing breakthrough performances in various protein-related tasks. However, substantial work is still required before these models can reliably perform highly specific tasks, such as predicting TCR specificity to novel epitopes. Importantly, alongside the rapid development of new models, targeted efforts to understand the strengths and weaknesses of each model need to be undertaken. In this direction, the ImmRep 2022 Workshop [9] has initiated efforts to create a common benchmark dataset and to conduct yearly competitions to facilitate rigorous model comparisons–a critical advancement for the field.
The interpretation of existing models is also necessary to improve our understanding of TCR-epitope interactions. Many machine learning models, especially PLMs, are inherently non-interpretable due to the encoding of amino acid information in highly abstract and complex latent spaces. The challenge of deriving meaning and explanations from black-box models presents a crucial challenge that could be addressed by leveraging recent advancements in Interpretable Machine Learning. For instance, the Automated Concept-based Explanation (ACE) [65] aims to automatically identify key concepts, i.e. high-level, human-understandable features, that influence model decisions. Applied to clusters of samples in the latent space, it could be used to explore the biochemical properties that govern cluster assignments.
Attention mechanisms can also be leveraged to capture structurally important residue pairs that contribute to TCR-epitope binding [66, 67] or predict protein structural properties [68]. For instance, a recent study categorized residues into groups with high and low attention values. The analysis revealed that residues receiving more attention often had distinct structural properties, and they were statistically more likely to form hydrogen bonds within the CDR3 region [68]. Alternatively, black-box models can be employed to generate plausible hypotheses that are later investigated using post hoc interpretable methods. An example of this approach is DECODE [69], a user-centric interpretable pipeline to extract human-comprehensible rules from any black-box TCR predictive model. DECODE exploits Anchors, a model-agnostic approach to approximate the decision boundary of any machine learning model and identify local, sufficient conditions for predictions [70]. However it is addressed, interpreting model predictions is not just a technical necessity but a fundamental requirement for clinical trust and application [71, 72, 73].
Other open questions demand nuanced investigation, such as understanding how PLMs organize information in the latent space. Counterintuitively, recent research has shown that the selecting of an appropriate PLM embedding layer for downstream applications necessitates optimization, as in some instances, earlier layers demonstrate superior representation capabilities for specific tasks [64]. Gaining a deeper insight into what and how information is encoded could enable more focused and efficient development and deployment of models for particular tasks.
Finally, many computational methods designed for TCR specificity prediction can be adapted to model B cell receptor (BCR) antigen binding. Indeed, both statistical [74], machine learning [75], and transformer-based approaches [76] have been developed to jointly model TCR and BCR repertoires. While not the primary focus of this paper, BCR specificity prediction faces similar data and computational challenges as TCR specificity prediction. A particularly interesting question is whether PLM-based models can effectively capture both TCR and BCR characteristics, or if specialized models might yield higher accuracy.
@@
In summary, the field of computational immunology, particularly regarding TCR and BCR specificity prediction, is rapidly evolving, yet key questions and challenges persist. Addressing these is crucial not only for deepening our understanding of the immune system but also for paving the way for groundbreaking clinical applications.
@@ 7 Appendix
Survey of TCR specificity prediction models.
@@
| diffRBM [86] | 2023 | Supervised | No |

### v2

@@ 5 Outlook
The accurate prediction of TCR specificity is crucial for various clinical applications, including the development of safer and more efficient immunotherapies and gaining a deeper understanding of autoimmune diseases. Despite the progress, current TCR prediction models face significant challenges due to insufficient and often biased data, notably concerning epitope information. This limitation restricts the models’ capacity to generalize to new epitopes, which is critical, for instance, to predict potential cross-reactive events in T cell-based therapies.
In recent years, the field of immunoinformatics has significantly benefited from the integration of machine learning techniques with proven efficacy in other domains, including established neural network architectures such as Convolutional Neural Networks (CNNs) and Recurrent Neural Networks (RNNs), transfer learning strategies, and NLP-inspired approaches such as word embeddings. The recent adoption of Transformers and language models has led to the emergence of protein language models (PLMs), which have substantially improved the accuracy of computational models for protein analysis and prediction, including TCR prediction models. This innovation is revolutionizing performance in data-scarce environments and demonstrating remarkable successes across various protein-centric tasks. Despite these advancements, considerable work remains to be done, particularly in enabling these models to reliably execute highly specific functions, such as predicting TCR specificity to novel epitopes.
[…]
Finally, many computational methods designed for TCR specificity prediction can be adapted to model B cell receptor (BCR) antigen binding. Indeed, both statistical [109], machine learning [110], and Transformer-based approaches [111] have been developed to jointly model TCR and BCR repertoires. While not the primary focus of this paper, BCR specificity prediction faces similar data and computational challenges as TCR specificity prediction. However, BCR-antigen binding typically involves non-linear epitopes, unlike the linear epitopes involved in T-cell peptide binding. Furthermore, BCR datasets often list the target antigen with an unknown epitope, making the BCR problem generally more challenging. Addressing these challenges might involve adapting TCR models for BCR modeling by modifying the algorithms to incorporate structural information about non-linear epitopes. Additionally, improving the resolution of antigen mapping, employing advanced algorithms capable of handling full antigen sequences, and leveraging transfer learning to fine-tune models for BCR-related tasks will also be beneficial. Ultimately, an intriguing question remains whether PLM-based models can effectively capture both TCR and BCR characteristics or if specialized models might yield higher accuracy. Similarly, the question of how PLMs organize information in the latent space requires deeper investigation. A better understanding of what and how information is encoded could lead to more focused and efficient development and deployment of models for both TCR and BCR-related tasks.
@@
In summary, the field of computational immunology, particularly in TCR specificity prediction, is rapidly evolving, yet key questions and challenges persist. Addressing these issues is crucial not only for deepening our understanding of the immune system but also for paving the way for groundbreaking clinical applications.
@@ 7 Appendix
Survey of TCR specificity prediction models (published by December 2023).
@@
| SwarmTCR [40] | 2021 | Supervised | No |
| ClusTCR [110] | 2021 | Unsupervised | No |
| SONIA [33] | 2021 | Supervised | No |
@@
| PiTE [118] | 2022 | Language | Yes |
| Bi et al [119] | 2022 | Supervised | Yes |
| diffRBM [120] | 2023 | Supervised | No2 |